# AXIS: Explainable Time Series Anomaly Detection with Large Language Models

### v1

@@ Abstract
Time-series anomaly detection (TSAD) increasingly demands explanations that articulate not only if an anomaly occurred, but also what pattern it exhibits and why it is anomalous. Leveraging the impressive explanatory capabilities of Large Language Models (LLMs), recent works have attempted to treat time series as text for explainable TSAD. However, this approach faces a fundamental challenge: LLMs operate on discrete tokens and struggle to directly process long, continuous signals. Consequently, naïve time-to-text serialization suffers from a lack of contextual grounding and representation alignment between the two modalities. To address this gap, we introduce AXIS, a framework that conditions a frozen LLM for nuanced time-series understanding. Instead of direct serialization, AXIS enriches the LLM’s input with three complementary hints derived from the series: (i) a symbolic numeric hint for numerical grounding, (ii) a context-integrated, step-aligned hint distilled from a pretrained time-series encoder to capture fine-grained dynamics, and (iii) a task-prior hint that encodes global anomaly characteristics. Furthermore, to facilitate robust evaluation of explainability, we introduce a new benchmark featuring multi-format questions and rationales that supervise contextual grounding and pattern-level semantics. Extensive experiments, including both LLM-based and human evaluations, demonstrate that AXIS yields explanations of significantly higher quality and achieves competitive detection accuracy compared to general-purpose LLMs, specialized time-series LLMs, and time-series Vision Language Models.
@@ 1 Introduction
Time Series Anomaly Detection (TSAD) is essential for safeguarding critical systems across domains (Iqbal et al., 2019; Zeufack et al., 2021; Hundman et al., 2018). While deep learning models can detect anomalies with high accuracy (Fig. 1(a)), their adoption in real-world systems is limited by two challenges. First, their reasoning process is essentially a black box. Experts remain in the dark when asking the most practical question: why was this anomaly flagged? Post-hoc attribution methods such as SHAP (Fig. 1(b)) attempt to fill this void, but they merely repackage correlations into statistical features. Such attributions reveal what inputs influenced the model, but they stop short of offering why the underlying anomaly event occurred. Second, these models are brittle. Trained narrowly - often on a single dataset - they capture dataset-specific features rather than generalizable patterns. In fast-changing environments where anomalies manifest in diverse forms, this rigidity is crippling: models must be retrained at high cost, yet still fail to transfer across domains. What is missing is the ability to handle anomalies universally—to recognize and adapt across diverse failure patterns without exhaustive retraining.
@@
In response to these limitations, the community has turned to Large Language Models (LLMs), celebrated for their fluency and generalization. Yet critical obstacles remain: LLMs operate on discrete tokens, making them ill-suited for the long, continuous nature of time series  (Dong et al., 2024). Attempting to fit these signals into tokenized inputs often incurs lossy serialization, forcing workarounds that undercut the LLM’s capabilities. Common strategies—feeding isolated fragments or pre-aggregated statistics—reduce the model to a mere post-hoc translator. As our motivating example illustrates (Fig. 2(a)), this naive approach, however, suffers a critical semantic gap, driven by two fundamental failures.
@@ 2 Related Work
Classical and statistical methods remain competitive baselines yet produce pointwise scores with weak semantics and limited support for multivariate structure (Liu et al., 2008; Yeh et al., 2016). Deep TSAD improves accuracy via reconstruction (Audibert et al., 2020; Su et al., 2019; Zhang et al., 2019), prediction-residual modeling (Tuli et al., 2022), and attention-centric architectures (Xu et al., 2021; Yang et al., 2023; Shen et al., 2020; Lan et al., 2025a), with recent work exploring unified/foundation-style formulations (Shentu et al., 2024; Gao et al., 2024) and diffusion-based detectors (Wang et al., 2025b). Explanations, however, are largely post hoc and tied to low-level contributions (e.g., time-step or feature importance), limiting mechanism-oriented diagnosis. Time-series XAI extends attribution (Bento et al., 2021) and investigates prototype/shapelet/motif views and counterfactual recourse (Bahri et al., 2022; Yeh et al., 2016), but explanations remain grounded in signal-level statistics rather than pattern-level concepts. This motivates treating TSAD explainability as a semantic alignment problem.
@@
LLMs can function as zero-shot anomaly detectors under appropriate prompting and input scaling (Alnegheimish et al., 2024; Dong et al., 2024; Zhou & Yu, 2024; Wang et al., 2025a). Performance degrades on long-horizon series due to context-length limits, lossy time-to-text serialization, and chunked inference, which together induce memory decay and boundary artifacts. A complementary line uses LLMs as post-hoc reasoners that verbalize anomaly scores, SHAP attributions, or raw subsequences, or coordinate multi-agent annotation (Liu et al., 2025; Lin et al., 2024). In both paradigms, LLMs act mainly as summarizers of low-level signals, yielding descriptive rather than semantically grounded explanations. Our approach directly aligns temporal representations with language via soft-prompt-based conditioning, aiming for faithful, pattern-level explanations.
@@
Existing benchmarks for time-series question answering, which are adjacent to our task, can be broadly categorized into two paradigms. The first relies on fully synthetic data generation, where normal time series are composed from trends, seasonality, and noise, after which localized anomalies are injected to generate templated or LLM-augmented labels (Cai et al., 2024; Xie et al., 2024; Wang et al., 2025a; Kong et al., 2025). The second paradigm uses real-world datasets, pairing authentic time-series data with corresponding semantic information to create evaluation suites (Kim et al., 2024; Cai et al., 2025; Liu et al., 2024a; Williams et al., 2024; Chen et al., 2025). However, synthetic benchmarks often lack the contextual richness required for robust grounding and representation alignment, while real-world data yields domain-specific explanations that limit model generalizability. To our knowledge, a dedicated benchmark for semantic time series anomaly explanation has remained a critical gap, which our work directly addresses by introducing a benchmark designed for this task.
@@ 3.2 AXIS Framework
LLMs possess native reasoning capabilities over discrete numerals when presented symbolically, even in zero-shot settings. To exploit this capability without exhausting the context budget, we textualize only the target window $[s,e)$ after Z-score normalization of the full series $\mathbf{x}_{1:T}$. Values are scaled by a factor $r$ (default r=100r{=}100) to preserve precision while avoiding decimal tokens (Liu et al., 2024b), rounded to integers, and serialized as a delimiter-separated string (e.g., “123, 124, 127, 128”) to constitute {time_series_tokens}. This pathway is compact—its position cost scales as $\alpha(e-s)+c$ where $\alpha$ is the average subword tokens per integer and $c$ the delimiter overhead—yet it preserves step-wise numeric grounding.
@@
While the above textualization provides direct numeric access, it cannot capture long-range dependencies essential for TSAD such as regime shifts, seasonality interactions, and boundary effects. We therefore condense global information into step-aligned local representations via a pretrained time-series encoder and a Hint Tuner, in the spirit of (Jin et al., 2023). A Transformer encoder $f_{\theta}$ consumes $\mathbf{x}_{1:T}$ and outputs embeddings $\mathbf{H}_{1:T}\in\mathbb{R}^{T\times d_{\text{proj}}}$ where $d_{\text{proj}}$ is the projection dimension (the details are shown in Appx. B.1). We slice $\mathbf{H}_{s:e}$ and map it into the LLM space using cross-attention over a prototype bank derived from the LLM vocabulary: $\mathbf{S}_{\text{proto}}=\mathbf{M}\mathbf{E}_{\text{vocab}}$ where $\mathbf{E}_{\text{vocab}}\in\mathbb{R}^{|\mathcal{V}|\times d_{h}}$ is fixed word embeddings and $\mathbf{M}\in\mathbb{R}^{P\times|\mathcal{V}|}$ is a learnable linear projection. Queries $\mathbf{W}_{q}\,\mathbf{H}_{s:e}$ attend to this bank: $\tilde{\mathbf{H}}_{s:e}=\mathrm{Attn}\!\big(\mathbf{W}_{q}\,\mathbf{H}_{s:e},\,\mathbf{S}_{\text{proto}},\,\mathbf{S}_{\text{proto}}\big)\ \in\ \mathbb{R}^{(e-s)\times d_{h}}.$ The resulting $\tilde{\mathbf{H}}_{s:e}$ acts as step-aligned local hints that inject global context and temporal structure into the LLM embedding space while keeping both the LLM and $f_{\theta}$ frozen. This pathway adds (es)(e-s) positions linearly while supplying detailed global context and temporal alignment.
@@ 3.4 A Benchmark for semantic time series anomaly explanation
To address the representation alignment challenge, we introduce a procedural engine that moves beyond abstract statistical deviations to a vocabulary of interpretable, pattern-level anomalies using the procedure of Lan et al. (2025b). As illustrated in Figure 4, our engine synthetically composes canonical anomaly primitives—such as sudden spikes, level shifts, and periodicity breaks—onto clean baseline series. A key advantage of our approach is the generation of paired time series: for every abnormal series created, a corresponding normal counterpart is preserved. This methodology establishes an unambiguous, verifiable link between anomaly time series and its linguistic label, forming the bedrock for teaching models to reason about the semantics of temporal events.
@@ 4.1 Experimental Setup
We compare AXIS with a comprehensive set of strong baselines, categorized as follows: Timeseries VLM:(He et al., 2025) Image LLM is supported by gpt-4o, which analyzes plots of the full time series with highlighted window, treating the explanation task as a visual reasoning problem. Specialized TS-LLM Methods: We include several recent models designed for time series analysis with LLMs: ChatTS (Xie et al., 2024), LLMAD (Liu et al., 2025), ChatTime (Wang et al., 2025a), and AnomLLM (Dong et al., 2024). We evaluate AnomLLM in two settings: providing the full series (AnomLLM(Full)) and providing only the target window (AnomLLM(Window)).
@@
Following recent work on evaluating LLM-generated content, we use an LLM-as-a-judge approach, specifically G-eval (Liu et al., 2023) with Gemini-2.5 as the arbiter. The quality of explanations is assessed across multiple dimensions tailored to each question type, including Correctness (Corr.), Reasoning Quality (Rsn. Qual.), Accuracy (Acc.), Completeness (Comp.), Relevance (Rel.), and Justification Quality (Justif.). A final, holistic score (Final) is also computed. The detailed definition for evaluation metrics is given in Appx. E
@@ Appendix D Phase I TSAD results
The comparison covers classical, deep generative, and foundation-style models. IForest (Liu et al., 2008) and LOF (Breunig et al., 2000) serve as nonparametric unsupervised baselines. OmniAnomaly (Su et al., 2019), USAD (Audibert et al., 2020), and TranAD (Tuli et al., 2022) represent deep unsupervised anomaly detectors. DADA (Shentu et al., 2024) and TSPulse (Ekambaram et al., 2025), MOMENT_ZS (Goswami et al., 2024), Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024), and Time_MOE (Xiaoming et al., 2025) are pretrained time-series models evaluated in zero-shot or lightly adapted settings. AXIS denotes a semantic representation-based detector. Implementations follow public releases or authors’ configurations when feasible.

### v2

@@ Abstract
Time-series anomaly detection (TSAD) increasingly demands explanations that articulate not only if an anomaly occurred, but also what pattern it exhibits and why it is anomalous. Leveraging the impressive explanatory capabilities of Large Language Models (LLMs), recent works have attempted to treat time series as text for explainable TSAD. However, this approach faces a fundamental challenge: LLMs operate on discrete tokens and struggle to directly process long, continuous signals. Consequently, naïve time-to-text serialization suffers from a lack of contextual grounding and representation alignment between the two modalities. To address this gap, we introduce AXIS, a framework that conditions a frozen LLM for nuanced time-series understanding. Instead of direct serialization, AXIS enriches the LLM’s input with three complementary hints derived from the series: (i) a symbolic numeric hint for numerical grounding, (ii) a context-integrated, step-aligned hint distilled from a pretrained time-series encoder to capture fine-grained dynamics, and (iii) a task-prior hint that encodes global anomaly characteristics. Furthermore, to facilitate robust evaluation of explainability, we introduce a new benchmark featuring multi-format questions and rationales that supervise contextual grounding and pattern-level semantics. Extensive experiments, including both LLM-based and human evaluations, demonstrate that AXIS yields explanations of significantly higher quality and achieves competitive detection accuracy compared to general-purpose LLMs, specialized time-series LLMs, and time-series Vision Language Models.  [^1] Code is available at https://github.com/thu-sail-lab/AXIS.
@@ 1 Introduction
Time Series Anomaly Detection (TSAD) is essential for safeguarding critical systems across domains (iqbal2019fault; zeufack2021unsupervised; hundman2018detecting). While deep learning models can detect anomalies with high accuracy (Fig. 1(a)), their adoption in real-world systems is limited by two challenges. First, their reasoning process is essentially a black box. Experts remain in the dark when asking the most practical question: why was this anomaly flagged? Post-hoc attribution methods such as SHAP (Fig. 1(b)) attempt to fill this void, but they merely repackage correlations into statistical features. Such attributions reveal what inputs influenced the model, but they stop short of offering why the underlying anomaly event occurred. Second, these models are brittle. Trained narrowly - often on a single dataset - they capture dataset-specific features rather than generalizable patterns. In fast-changing environments where anomalies manifest in diverse forms, this rigidity is crippling: models must be retrained at high cost, yet still fail to transfer across domains. What is missing is the ability to handle anomalies universally—to recognize and adapt across diverse failure patterns without exhaustive retraining.
@@
In response to these limitations, the community has turned to Large Language Models (LLMs), celebrated for their fluency and generalization. Yet critical obstacles remain: LLMs operate on discrete tokens, making them ill-suited for the long, continuous nature of time series  (dong2024can). Attempting to fit these signals into tokenized inputs often incurs lossy serialization, forcing workarounds that undercut the LLM’s capabilities. Common strategies—feeding isolated fragments or pre-aggregated statistics—reduce the model to a mere post-hoc translator. As our motivating example illustrates (Fig. 2(a)), this naive approach, however, suffers a critical semantic gap, driven by two fundamental failures.
@@ 2 Related Work
Classical and statistical methods remain competitive baselines yet produce pointwise scores with weak semantics and limited support for multivariate structure (liu2008isolation; yeh2016matrix). Deep TSAD improves accuracy via reconstruction (audibert2020usad; su2019robust; zhang2019deep), prediction-residual modeling (tuli2022tranad), and attention-centric architectures (xu2021anomaly; yang2023dcdetector; shen2020timeseries; lan2025cicada), with recent work exploring unified/foundation-style formulations (shentu2024towards; gao2024units) and diffusion-based detectors (wang2025diffad). Explanations, however, are largely post hoc and tied to low-level contributions (e.g., time-step or feature importance), limiting mechanism-oriented diagnosis. Time-series XAI extends attribution (bento2021timeshap) and investigates prototype/shapelet/motif views and counterfactual recourse (bahri2022shapelet; yeh2016matrix), but explanations remain grounded in signal-level statistics rather than pattern-level concepts. This motivates treating TSAD explainability as a semantic alignment problem.
@@
LLMs can function as zero-shot anomaly detectors under appropriate prompting and input scaling (alnegheimish2024can; dong2024can; zhou2024can; wang2025chattime). Performance degrades on long-horizon series due to context-length limits, lossy time-to-text serialization, and chunked inference, which together induce memory decay and boundary artifacts. A complementary line uses LLMs as post-hoc reasoners that verbalize anomaly scores, SHAP attributions, or raw subsequences, or coordinate multi-agent annotation (liu2025large; lin2024decoding). In both paradigms, LLMs act mainly as summarizers of low-level signals, yielding descriptive rather than semantically grounded explanations. Our approach directly aligns temporal representations with language via soft-prompt-based conditioning, aiming for faithful, pattern-level explanations.
@@
Existing benchmarks for time-series question answering, which are adjacent to our task, can be broadly categorized into two paradigms. The first relies on fully synthetic data generation, where normal time series are composed from trends, seasonality, and noise, after which localized anomalies are injected to generate templated or LLM-augmented labels (cai2024timeseriesexam; xie2024chatts; wang2025chattime; kong2025time2). The second paradigm uses real-world datasets, pairing authentic time-series data with corresponding semantic information to create evaluation suites (kim2024multi; cai2025timeseriesgym; liu2024time2; williams2024context; chen2025mtbench). However, synthetic benchmarks often lack the contextual richness required for robust grounding and representation alignment, while real-world data yields domain-specific explanations that limit model generalizability. To our knowledge, a dedicated benchmark for semantic time series anomaly explanation has remained a critical gap, which our work directly addresses by introducing a benchmark designed for this task.
@@ 3.2 AXIS Framework
LLMs possess native reasoning capabilities over discrete numerals when presented symbolically, even in zero-shot settings. To exploit this capability without exhausting the context budget, we textualize only the target window $[s,e)$ after Z-score normalization of the full series $\mathbf{x}_{1:T}$. Values are scaled by a factor $r$ (default r=100r{=}100) to preserve precision while avoiding decimal tokens (liu2024time), rounded to integers, and serialized as a delimiter-separated string (e.g., “123, 124, 127, 128”) to constitute {time_series_tokens}. This pathway is compact—its position cost scales as $\alpha(e-s)+c$ where $\alpha$ is the average subword tokens per integer and $c$ the delimiter overhead—yet it preserves step-wise numeric grounding.
@@
While the above textualization provides direct numeric access, it cannot capture long-range dependencies essential for TSAD such as regime shifts, seasonality interactions, and boundary effects. We therefore condense global information into step-aligned local representations via a pretrained time-series encoder and a Hint Tuner, in the spirit of (jin2023time). A Transformer encoder $f_{\theta}$ consumes $\mathbf{x}_{1:T}$ and outputs embeddings $\mathbf{H}_{1:T}\in\mathbb{R}^{T\times d_{\text{proj}}}$ where $d_{\text{proj}}$ is the projection dimension (the details are shown in Appx. B.1). We slice $\mathbf{H}_{s:e}$ and map it into the LLM space using cross-attention over a prototype bank derived from the LLM vocabulary: $\mathbf{S}_{\text{proto}}=\mathbf{M}\mathbf{E}_{\text{vocab}}$ where $\mathbf{E}_{\text{vocab}}\in\mathbb{R}^{|\mathcal{V}|\times d_{h}}$ is fixed word embeddings and $\mathbf{M}\in\mathbb{R}^{P\times|\mathcal{V}|}$ is a learnable linear projection. Queries $\mathbf{W}_{q}\,\mathbf{H}_{s:e}$ attend to this bank: $\tilde{\mathbf{H}}_{s:e}=\mathrm{Attn}\!\big(\mathbf{W}_{q}\,\mathbf{H}_{s:e},\,\mathbf{S}_{\text{proto}},\,\mathbf{S}_{\text{proto}}\big)\ \in\ \mathbb{R}^{(e-s)\times d_{h}}.$ The resulting $\tilde{\mathbf{H}}_{s:e}$ acts as step-aligned local hints that inject global context and temporal structure into the LLM embedding space while keeping both the LLM and $f_{\theta}$ frozen. This pathway adds (es)(e-s) positions linearly while supplying detailed global context and temporal alignment.
@@ 3.4 A Benchmark for semantic time series anomaly explanation
To address the representation alignment challenge, we introduce a procedural engine that moves beyond abstract statistical deviations to a vocabulary of interpretable, pattern-level anomalies using the procedure of lan2025towards. As illustrated in Figure 4, our engine synthetically composes canonical anomaly primitives—such as sudden spikes, level shifts, and periodicity breaks—onto clean baseline series. A key advantage of our approach is the generation of paired time series: for every abnormal series created, a corresponding normal counterpart is preserved. This methodology establishes an unambiguous, verifiable link between anomaly time series and its linguistic label, forming the bedrock for teaching models to reason about the semantics of temporal events.
@@ 4.1 Experimental Setup
We compare AXIS with a comprehensive set of strong baselines, categorized as follows: Timeseries VLM:(he2025harnessing) Image LLM is supported by gpt-4o, which analyzes plots of the full time series with highlighted window, treating the explanation task as a visual reasoning problem. Specialized TS-LLM Methods: We include several recent models designed for time series analysis with LLMs: ChatTS (xie2024chatts), LLMAD (liu2025large), ChatTime (wang2025chattime), and AnomLLM (dong2024can). We evaluate AnomLLM in two settings: providing the full series (AnomLLM(Full)) and providing only the target window (AnomLLM(Window)).
@@
Following recent work on evaluating LLM-generated content, we use an LLM-as-a-judge approach, specifically G-eval (liu2023g) with Gemini-2.5 as the arbiter. The quality of explanations is assessed across multiple dimensions tailored to each question type, including Correctness (Corr.), Reasoning Quality (Rsn. Qual.), Accuracy (Acc.), Completeness (Comp.), Relevance (Rel.), and Justification Quality (Justif.). A final, holistic score (Final) is also computed. The detailed definition for evaluation metrics is given in Appx. E
@@ Appendix D Phase I TSAD results
The comparison covers classical, deep generative, and foundation-style models. IForest (liu2008isolation) and LOF (breunig2000lof) serve as nonparametric unsupervised baselines. OmniAnomaly (su2019robust), USAD (audibert2020usad), and TranAD (tuli2022tranad) represent deep unsupervised anomaly detectors. DADA (shentu2024towards) and TSPulse (ekambaram2025tspulse), MOMENT_ZS (goswami2024moment), Chronos (ansari2024chronos), TimesFM (das2024decoder), and Time_MOE (xiaoming2025time) are pretrained time-series models evaluated in zero-shot or lightly adapted settings. AXIS denotes a semantic representation-based detector. Implementations follow public releases or authors’ configurations when feasible.